### v1

@@ Abstract
Non-invasive Brain-Computer Interfaces (BCI) offer a safe and accessible means of connecting the human brain to external devices, with broad applications in home and clinical settings to enhance human capabilities. However, the high noise level and limited task-specific data in non-invasive signals constrain decoding capabilities. Recently, the adoption of self-supervised pre-training is transforming the landscape of non-invasive BCI research, enabling the development of brain foundation models to capture generic neural representations from large-scale unlabeled electroencephalography (EEG) signals with substantial noises. However, despite these advances, the field currently lacks comprehensive, practical and extensible benchmarks to assess the utility of the public foundation models across diverse BCI tasks, hindering their widespread adoption. To address this challenge, we present AdaBrain-Bench, a large-scale standardized benchmark to systematically evaluate brain foundation models in widespread non-invasive BCI tasks. AdaBrain-Bench encompasses a diverse collection of representative BCI decoding datasets spanning 7 key applications. It introduces a streamlined task adaptation pipeline integrated with multi-dimensional evaluation metrics and a set of adaptation tools. The benchmark delivers an inclusive framework for assessing generalizability of brain foundation models across key transfer settings, including cross-subject, multi-subject, and few-shot scenarios. We leverage AdaBrain-Bench to evaluate a suite of publicly available brain foundation models and offer insights into practices for selecting appropriate models in various scenarios. We make our benchmark pipeline available at Github repository to enable reproducible research and external use, offering a continuously evolving platform to foster progress toward robust and generalized neural decoding solutions.
@@ 1 Introduction
Figure 1: Overview of AdaBrain-Bench. a, categorization of 13 benchmark datasets and 7 benchmark tasks ranging from cognitive state assessment, human augmentation, and clinical monitoring. b, diverse task paradigms including classification, regression, and retrieval paired with comprehensive evaluation metrics. c, evaluated baselines including several recently open-sourced self-supervised foundation models (BIOT [21], EEGPT [22], LaBraM [23], and CBraMod [24]) and supervised traditional models (EEGNet [25], LDMA [26], ST-Tran [27], and Conformer [28]) used for comparison. d, multifaceted evaluation settings including cross-subject transfer, multi-subject adaptation and few-shot transfer settings to thoroughly assess models downstream task generalization ability in various scenarios.
@@ 4 Methods
Cognitive State Assessment Emotion Recognition SEED [46] 1, [^0] Hz 62 1s 15 144,852 B-Acc F1-W
@@
Workload Classification EEGMAT [48, 49] 500Hz 19 4s 36 1,080 B-Acc AUROC
Vigilance Estimation SEED-VIG [50] 200Hz 17 8s 21 20,355 Corr R2
Human Augmentation Motor Imagery Classification BCI-IV-2A [51] 250Hz 22 4s 9 5,184 B-Acc F1-W
@@
Visual Decoding Things-EEG [53, 54] 1, [^0] Hz 63 1s 10 821,600 2-Way Top-5
Clinical Monitoring Clinical Anomaly Detection TUEV [56] 256Hz 23 5s 370 112,237 B-Acc F1-W
@@
Normalization. Data normalization is essential to accommodate varying data distributions across EEG datasets. We applied z-score normalization as the standard technique for most datasets, which normalizes each sample using global statistics across all trials to conform to a normal distribution. For SHU dataset, which is prone to outliers, we used 95-percentile normalization to enhance robustness, which normalizes data points by dividing them by the 95-th percentile of absolute amplitude per channel in each trial, thus mitigating the impact of outliers. For Things-EEG dataset, we directly rescaled the signal to ensure that most signal values fall within the range of approximately $[-1,1]$ following previous practice [23].
@@
Other Experimental Details. We use binary cross-entropy (BCE) loss for binary classification, cross-entropy (CE) loss for multiclass classification, mean squared error (MSE) for regression, and contrastive loss for retrieval. We perform a learning rate grid search over $[1e-3,5e-4,1e-4,5e-4,1e-5]$ for each adaptation strategy/setting combination to determine the optimal value. The batch size is set as 64. We adopted the AdamW optimizer, with the weight decay set as 0.05. The number of training epoch is set as 50. The experiments are conducted using Python 3.10.13 and PyTorch 2.4.0 with CUDA 11.8 support. Notably, for foundation models (e.g., BIOT and EEGPT) unable to directly decode downstream datasets due to channel system incompatibilities, we applied a $1\times 1$ channel-wise convolutional layer to ensure compatibility with input requirements.
@@
Accuracy Metrics. For binary classification tasks, different evaluation metrics are adopted based on the characteristics of each dataset. For balanced binary classification tasks (e.g., EEGMAT and SHU), we report the Balanced Accuracy (B-Acc) and the Area Under the Receiver Operating Characteristic Curve (AUROC). B-Acc is defined as the average recall across all classes, while AUROC quantifies the model’s overall discriminative ability across different classification thresholds. For imbalanced classification tasks (e.g., TUAB and Siena), where the positive class is of particular importance, we report B-Acc and the Area Under the Precision-Recall Curve (AUC-PR), as AUC-PR more effectively reflects the model’s performance on the minority (positive) class. For multiclass classification tasks (i.e., SEED, SEED-IV, EEGMAT, BCI-IV-2A, TUEV, HMC, SHHS, and Sleep-EDF), we use B-Acc and the Weighted F1-Score (F1-W) as the metric. F1-W evaluates model performance by calculating a weighted average of each class’s F1 score, where weights are determined by class sample proportion to alleviate class imbalance issue. For regression tasks (i.e., SEED-VIG), we adopt Pearson Correlation Coefficient (Corr) and R2 score as the metric. Corr measures the linear relationship between predicted and true values, while R2 represents the proportion of variance in the dependent variable explained by the decoding model. For retrieval tasks (i.e., Things-EEG), we use 2-Way and Top-5 accuracy as the metric, following the previous practice [55]. The N-way accuracy is computed by constructing a set containing the target sample and $N-1$ distractor samples from the test set, assessing the model’s ability to identify the correct target among $N$ categories. Top-5 accuracy measures the frequency that the true target image is among the top 5 retrieved results, based on the corresponding EEG signals.
@@
Model #Params. Architecture Pretraining Strategy Pretraining Data Size (#channels * h) Pretraining Signal Duration Pre-training Datasets #Supported Channels
BIOT [21] 3.2M Linear Transformer Self-supervised Contrastive Learning 312,044 60,682h PREST, SHHS, CHB-MIT, IIIC Seizure, TUAB & TUEV 18
EEGPT [22] 25M CNN & Transformer Masked Signal Reconstruction 16,550 246h PhysioNet-MI, HGD, TSU, SEED & M3CV 62
LaBraM [23] 5.8M CNN & Transformer Masked Signal Reconstruction 76,801 2,538h BCI-IV-1, Emobrain, SPIS, Grasp and Lift EEG, Inria P300, SEED series, Siena, TUEP, TUSZ & others 137
CBraMod [24] 4.0M CNN & Transformer Masked Signal Reconstruction 514,178 27,062h TUAB, TUAR TUEP, TUEV TUSE & TUSL Flexible
@@
EEGPT [22]. EEGPT is a transformer-based foundation model adopting mask-based self-supervised learning. The pretraining involves a mask-based reconstruction strategy and a spatio-temporal representation alignment strategy to learn generic neural representations from unlabeled EEG data. EEGPT was pretrained on 246 hours of EEG signals from five datasets from various tasks, rendering a large-scale foundation model with 25 million parameters and 62 supported channels.
@@
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory.

### v2

@@ Abstract
Non-invasive Brain-Computer Interfaces (BCI) offer a safe and accessible means of connecting the human brain to external devices, with broad applications in home and clinical settings to enhance human capabilities. However, the high noise level and limited task-specific data in non-invasive signals constrain decoding capabilities. Recently, the adoption of self-supervised pre-training is transforming the landscape of non-invasive BCI research, enabling the development of brain foundation models to capture robust neural representations from large-scale unlabeled electroencephalography (EEG) signals. However, despite these advances, the field currently lacks standard and comprehensive benchmarks to assess the utility of the public foundation models across diverse BCI tasks, hindering their widespread adoption. To address this challenge, we present AdaBrain-Bench, a large-scale standardized benchmark to systematically evaluate brain foundation models in widespread non-invasive BCI tasks. AdaBrain-Bench encompasses a diverse collection of representative BCI decoding datasets spanning 7 key applications. It introduces a streamlined model adaptation pipeline integrated with multi-dimensional evaluation metrics and a set of adaptation tools for flexible deployment. The benchmark establishes multiple transfer settings, including cross-subject, multi-subject, and few-shot settings, to rigorously evaluate models’ transferability across diverse real-world application scenarios. We leverage AdaBrain-Bench to evaluate publicly available brain foundation models and offer insights into practices for selecting appropriate models in various scenarios. We make our benchmark pipeline available at GitHub repository to enable reproducible research and external use, offering a continuously evolving platform to foster progress toward robust and generalized neural decoding solutions.
@@ 1 Introduction
Figure 1: Overview of AdaBrain-Bench. a, categorization of 13 benchmark datasets and 7 benchmark tasks ranging from cognitive state assessment, human augmentation, and clinical monitoring. b, diverse task paradigms including classification, regression, and retrieval paired with comprehensive evaluation metrics. c, evaluated baselines including several recently open-sourced self-supervised foundation models (BIOT [21], EEGPT [22], LaBraM [23], and CBraMod [24]) and supervised traditional models (EEGNet [25], LDMA [26], ST-Tran [27], and Conformer [28]) used for comparison. d, multifaceted evaluation settings including cross-subject transfer, multi-subject adaptation and few-shot transfer settings to thoroughly assess models downstream task generalization ability in various scenarios. e, the overall foundation model adaptation pipeline. Users input their model to the benchmark and brain signal data from their selected task.
@@ 4 Methods
Cognitive State Assessment
Emotion
Recognition
SEED [46] 1, [^0] Hz 62 1s 15 144,852 B-Acc F1-W
@@
Workload
Classification
EEGMAT [48, 49] 500Hz 19 4s 36 1,080 B-Acc AUROC
Vigilance
Estimation
SEED-VIG [50] 200Hz 17 8s 21 20,355 Corr R2
Human Augmentation
Motor Imagery
Classification
BCI-IV-2A [51] 250Hz 22 4s 9 5,184 B-Acc F1-W
@@
Visual
Decoding
Things-EEG [53, 54] 1, [^0] Hz 63 1s 10 821,600 2-Way Top-5
Clinical Monitoring
Clinical Anomaly
Detection
TUEV [56] 256Hz 23 5s 370 112,237 B-Acc F1-W
@@
Normalization. Data normalization is essential to accommodate varying data distributions across EEG datasets. We applied z-score normalization as the standard technique for most datasets, which normalizes each sample using global statistics across all trials to conform to a normal distribution. For SHU dataset, which is prone to outliers, we used 95-percentile normalization to enhance robustness, which normalizes data points by dividing them by the 95-th percentile of absolute amplitude per channel in each trial, thus mitigating the impact of outliers. For Things-EEG dataset, we directly rescaled the signal to ensure that most signal values fall within the range of approximately [1,1][-1,1] following previous practice [23].
@@
Other Experimental Details. We use binary cross-entropy (BCE) loss for binary classification, cross-entropy (CE) loss for multiclass classification, mean squared error (MSE) for regression, and contrastive loss for retrieval. We perform a learning rate grid search over [1e3,5e4,1e4,5e4,1e5][1e-3,5e-4,1e-4,5e-4,1e-5] for each adaptation strategy/setting combination to determine the optimal value. The batch size is set as 64. We adopted the AdamW optimizer, with the weight decay set as 0.05. The number of training epoch is set as 50. The experiments are conducted using Python 3.10.13 and PyTorch 2.4.0 with CUDA 11.8 support. Notably, for foundation models (e.g., BIOT and EEGPT) unable to directly decode downstream datasets due to channel system incompatibilities, we applied a $1\times 1$ channel-wise convolutional layer to ensure compatibility with input requirements.
@@
Accuracy Metrics. For binary classification tasks, different evaluation metrics are adopted based on the characteristics of each dataset. For balanced binary classification tasks (e.g., EEGMAT and SHU), we report the Balanced Accuracy (B-Acc) and the Area Under the Receiver Operating Characteristic Curve (AUROC). B-Acc is defined as the average recall across all classes, while AUROC quantifies the model’s overall discriminative ability across different classification thresholds. For imbalanced classification tasks (e.g., TUAB and Siena), where the positive class is of particular importance, we report B-Acc and the Area Under the Precision-Recall Curve (AUC-PR), as AUC-PR more effectively reflects the model’s performance on the minority (positive) class. For multiclass classification tasks (i.e., SEED, SEED-IV, EEGMAT, BCI-IV-2A, TUEV, HMC, SHHS, and Sleep-EDF), we use B-Acc and the Weighted F1-Score (F1-W) as the metric. F1-W evaluates model performance by calculating a weighted average of each class’s F1 score, where weights are determined by class sample proportion to alleviate class imbalance issue. For regression tasks (i.e., SEED-VIG), we adopt Pearson Correlation Coefficient (Corr) and R2 score as the metric. Corr measures the linear relationship between predicted and true values, while R2 represents the proportion of variance in the dependent variable explained by the decoding model. For retrieval tasks (i.e., Things-EEG), we use 2-Way and Top-5 accuracy as the metric, following the previous practice [55]. The N-way accuracy is computed by constructing a set containing the target sample and N1N-1 distractor samples from the test set, assessing the model’s ability to identify the correct target among $N$ categories. Top-5 accuracy measures the frequency that the true target image is among the top 5 retrieved results, based on the corresponding EEG signals.
@@
Model #Params. Architecture
Pretraining
Strategy
Pretraining Data Size
(#channels * h)
Pretraining
Signal Duration
Pre-training Datasets
#Supported
Channels
BIOT [21] 3.2M Linear Transformer
Self-supervised
Contrastive Learning
354,625 59,304h
PREST, SHHS, CHB-MIT,
IIIC Seizure, TUAB & TUEV
18
EEGPT [22] 25M CNN & Transformer
Masked Signal
Reconstruction
11,476 198h
PhysioNet-MI, HGD, TSU,
SEED & M3CV
58
LaBraM [23] 5.8M CNN & Transformer
Masked Signal
Reconstruction
76,801 2,535h
BCI-IV-1, Emobrain, SPIS,
Grasp and Lift EEG, Inria P300,
SEED series, Siena,
TUEP, TUSZ & others
136
CBraMod [24] 4.0M CNN & Transformer
Masked Signal
Reconstruction
175,678 9,246h
TUAB, TUAR
TUEP, TUEV
TUSE & TUSL
Flexible
@@
EEGPT [22]. EEGPT is a transformer-based foundation model adopting mask-based self-supervised learning. The pretraining involves a mask-based reconstruction strategy and a spatio-temporal representation alignment strategy to learn generic neural representations from unlabeled EEG data. EEGPT was pretrained on 246 hours of EEG signals from five datasets from various tasks, rendering a large-scale foundation model with 25 million parameters and 58 supported channels.